Prev Next

SW&IT Meeting Minutes 2022-06-20

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220613
LR20220613 - Bec, Matthieu 40.0
LR20220613 - Mechtley, Brandon 0.0
LR20220613 - Eychaner, Glenn 40.0
LR20220613 - Edmeades, David 40.0
LR20220613 - Grenz, Paul 36.0
LR20220613 - Hooper, Stephen 0.0
LR20220613 - Zhang, Yang 40.0
LR20220613 - Zhang, Xianyu 56.0

News from management

  • SHARK-NIR visit and arrival

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • UPS Tucson
  • Lasair III RMA

Team reports

MB

  • Last week
    • OC chair, ATT, SOUL-comm RIXes
    • APU AO stats scripts review/improvements
    • shark-vis PAE doc review
    • power outages support, TSS remote switch plans
    • Lasair III RMA/PO
  • This week
    • SOUL verification matrix meeting
    • shark-vis PAE doc review
    • vacation/remote Jun 25-Jul-31

YZ

  • Previous:
    • TMS test on Monday (has some problems).
    • Some TMS software work and extensive data analysis work.
    • Some TCS and misc support work due to power problems.
    • LBC and MODS misc tasks and code study and planning.
  • Next:
    • Continue to focus on TMS data analysis and debugging.
    • TMS software light refactoring to complete.
    • Other misc MODS/LBC and support work.

DE

  • Last week:
    • JS software, super annoying dance with Adobe but got there eventually.
      • Fun fact: screen saver will kick in on VNC logins and take all the CPU. I'll disable it for the LBTO user.
    • SE rsync: I think it's his DNS settings on the LBTI machines, but he didn't get back to me. I was able to successfully test from my workstation in Steward.
    • Begin particle monitor RMA process, LeRoy is building a <$250 shipping crate for it.
    • Obs3 disk unhappy? Will followup with LeRoy
  • This week:
    • Obs3
    • Tucson old Synology disk swap
    • Sort unhappy power out in Tucson server room
    • Coordinate particle monitor RMA with LeRoy as he has time to progress on it
    • Sanity check Veeam license limits before our per-socket licensing becomes unavailable.

XZ

  • Last week
    • Support shutdown and restart of DX and SX ASMs for 480 V non-UPS power removal
    • AO/SW meeting and discussions
    • Discussion and software change about fsm_skip_recovery
    • Review of AO night reports
    • support to LBC TMS test
    • handling of power loss on the mountain
    • planning for ATT activities
  • This week
    • DX ASM flat characterization
    • DX ASM MMC actuator list
    • SX ASM summer work preparations
    • AO support
    • Vacation (Jun 24 to Jul 8)

GE

  • Last week:
    • A new version of the RTD was installed in luci_rtd_test. It will likely be installed as the default next week. See RTD Troubleshooting and Problem Reporting for an overview.
    • IT 8355 (Aladin internal deadlocks): Closed. See IssueTrak for complete details.
    • IT 5815 (LUCI GUI slowdowns): Closed. See IssueTrak for complete details.
    • Worked with Olga and Dave T. on the N30 camera pinhole alignment process. The software is OK, but the boxes need to be recalibrated by Dave.
  • This week:
    • I will be unavailable all day June 20th.
    • I will be working half days (mornings 8-noon) most days through July 20th.
    • IT 8571, 8529 (Observer Scheduler GUI).

PSG

* Last week
    • LBTI SX/DX WFS machine fixes - CPU usage is now normal.
    • IT 8207: PCS printing error in tcslog (looks like it is merely a formatting error)
    • IT 4407: PCS: display the integral of the guiding corrections (Add numbers to PCSGUI)
    • IT 5357: add the HBS oil temperature to the alarm handler
    • power outages support, LBTI shutdown.
    • Testing new upgraded lbti-lmircam memory
  • This week
    • Continue work on issuetraks.

BM

  • Last week
    • Upgraded/switched to ao-supervisor stable on SX adsec/WFS, lbti-dxwfs
    • Emailed Fabio re: soul-dxwfs (badOcam branch + uncommitted changes), lbti-dxwfs (uncommitted changes)
    • Upgrade mishap on dxadsec, discussed procedure w/MB,XZ
    • Housekeeper: discussed w/XZ and made some changes (in PR), tested SBCU timeout in VM.
    • Sent out ao-supervisor status on adsec/WFS machines to aoops.
  • This week:
    • Pinged Alfio re: Q&A this morning, waiting to hear from Fabio re: DX WFS machines
    • Test plans for stable-mmc, Housekeeper, disable NCPA during optimize gain
    • ITs:
      • OffsetXY problems (RTDB disconnect in AOG meeting notes + IT 8307)
      • 8553 - bcu timeouts + ocam not live (8250/3957)
      • 8397 - aos disconnect

SH

  • Power-ups and downs on the mountain
  • Power issue in Tucson on Thursday
  • APMON

next:

  • Tucson CR PDU and work
  • hardware up the mountain
  • Room 507 computer upgrade
Topic revision: r10 - 20 Jun 2022, XianyuZhang
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback